OpenStudy (anonymous):

I've been struggling with this identity problem Establish the identity 1-1/2sin(2theta) = sin^3theta+cos^3theta/sintheta+costheta

OpenStudy (campbell_st):

well is the question 1. \[\frac{1}{2} \sin(2 \theta)\] or 2. \[\frac{1}{2\sin(2 \theta)}\]

OpenStudy (campbell_st):

oops forgot the 1 -

OpenStudy (anonymous):

The question is one minus one-half sin(2theta) equals sin cubed theta plus cos cubed theta over sine theta plus cos theta. So like [1-1/2\sin(2\theta)=\sin^3+\cos^3/\sin \theta+\cos\]

OpenStudy (anonymous):